# Impact of Dataset Properties on Membership Inference Vulnerability of Deep Transfer Learning

### v4

@@ 4.3 Model to predict dataset vulnerability
The trends seen in Figure 1 suggest the same power law relationship that we derived for the simplified model of membership inference in Section 3. We fit a linear regression model to predict $\log(\textsc{tpr}-\textsc{fpr})$ for each $\textsc{fpr}=10^{-k},k=1,\dots,5$ separately using the $\log C$ and $\log S$ as covariates with statsmodels (Seabold and Perktold, 2010). The general form of the model can be found in Equation 13, where $\beta_{S},\beta_{C}$ and $\beta_{0}$ are the learnable regression parameters.
@@
[…]
We utilise MIA results of ViT-B (Head) (see Table A3) as the training data. Based on the $R^{2}$ (coefficient of determination) score ($R^{2}=0.930$ for the model trained on $\textsc{fpr}=0.001$ data), our model fits the data extremely well. We provide further evidence for other fpr in Figure A.3 and Table A8 in the Appendix. Figure 4 shows the parameters of the prediction model fitted to the training data. For larger fpr, the coefficient $\beta_{S}$ is around $-0.5$, as our theoretical analysis predicts.
Figure 4: Coefficient values for different fpr when fitting a regression model based on Equation 13 fitted on data from ViT-B (Head) with LiRA (Table A3). The error bars display the 95% confidence intervals based on Student’s t-distribution. Theoretical values in the simplified model is shown by pink dotted lines ($\beta_{S}=0.5$ and $\beta_{C}=0$).
@@
  • R-50 (Head): Figure 5(a) shows that the regression model is robust to a change of the feature extractor, as it is able to predict the tpr for R-50 (Head) (test $R^{2}=0.790$).
  • R-50 (FiLM): Figure 5(b) shows that the prediction quality is good for R-50 (FiLM) models. These models are fine-tuned with parameter-efficient FiLM (Perez et al., 2018) layers (See Section C.1). Tobaben et al. (2023) demonstrated that FiLM layers are a competitive alternative to training all parameters. We supplement the MIA results of Tobaben et al. (2023) with own FiLM training runs. Refer to Table A7 in the Appendix.
  • From-Scratch-Training: Carlini et al. (2022) provide limited results on from-scratch-training. To the best of our knowledge these are the only published LiRA results on image classification models. Figure 5(b) displays that our prediction model underestimates the vulnerability of the from-scratch trained target models. We have identified two potential explanations for this 1 In from-scratch-training all weights of the model need to be trained from the sensitive data and thus potentially from-scratch-training could be more vulnerable than fine-tuning. 2 The strongest attack in Carlini et al. (2022) uses data augmentations to improve the performance. We are not using this optimization.
(a) The dots show the median tpr for the train set (ViT-B; Table A3) and the test set (R-50; Table A4) over six seeds (datasets: Patch Camelyon, EuroSAT and CIFAR100). The linear model is robust to changing the feature extractor from ViT-B to R-50.
(b) Regression model is robust to changing the fine-tuning method from Head to FiLM, but from scratch training seems to be more vulnerable than predicted. 1 left: fine-tuned with FiLM (see Table A7) 2 right: trained from scratch. Data is from Carlini et al. (2022) .
Figure 5: Performance of the regression model based on Equation 13 fitted on data from Table A3.
@@ 4.4 Individual MIA vulnerability
In order to assess the per-sample MIA vulnerability, we run the experiment with $257$ models with each of them once acting as a target and as a shadow model otherwise, compute the tpr at fpr for every sample separately. In Figure 6, we display the individual vulnerability as a function of $S$ (shots) for Patch Camelyon. The plot shows the maximal vulnerability of all samples and different quantiles over six seeds. The solid line is the median and the errorbars display the min and max over seeds. The quantiles are more robust to extreme outliers and show decreasing trends already at much lower $S$ than the maximum vulnerability.
When fitting the model in Equation 13 we observe that the coefficients $\beta_{S}$ that model the relationship between vulnerability and examples per class for the quantiles are $-0.5603$, $-0.5688$ and $-0.4796$ which is close to the theoretical value of $-0.5$ derived in the theoretical analysis in Section 3.4. However, the maximum vulnerability decreases with a lower slope of $-0.2695$ which is considerably smaller. With larger $S$ the slope of the max vulnerability increases, e.g., with $S\geq 32768$ the coefficient is $-0.3478$ suggesting that higher $S$ is required for the most vulnerable points.
@@
Table 1: The minimum $S$ with $C$=$2$ predicted by the models in Sections 4.3 and 4.4 to empirically match the DP bounds ($\delta$=$10^{-5}$) calculated through (Kairouz et al., 2015) in terms of tpr at fpr.
$\epsilon$ min $S$ in average case min $S$ in worst-case fpr=$0.1$ fpr=$0.01$ fpr=$0.001$ fpr=$0.1$ $0.25$ 5 40054005\,4005 400 69 0006900069\,00069 000 320 000320000320\,000320 000 $5.5\times 10^{9}$ $0.50$ 1 10011001\,1001 100 16 0001600016\,00016 000 88 0008800088\,00088 000 $2.6\times 10^{8}$ $0.75$ $360$ 5 90059005\,9005 900 38 0003800038\,00038 000 $3.5\times 10^{7}$ $1.00$ $160$ 2 70027002\,7002 700 19 0001900019\,00019 000 $7.0\times 10^{6}$
@@ 4.5 Comparison between empirical models and universal DP bounds
[…]
@@ 5 Discussion
Trying to bridge empirical MIA vulnerability and formal DP guarantees is not an easy task because of different threat models and different nature of bounds (statistical vs. universal). While we were able to show both theoretically (Section 3) and empirically (Section 4) that having more examples per class provides protection against MIA in fine-tuned neural networks, the numbers required for significant protection (Section 4.5) limit the practical utility of this observation.
@@
We mostly consider LiRA in our paper which is optimal for our simiplified model of membership inference (Section 3.4) but for the transfer-learning experiments (Section 4) there might be stronger attacks in the future. Furthermore, our results assume well-behaved underlying distributions. Formal bounds on MIA vulnerability would require something like DP. In addition, both our theoretical and empirical analysis focus on deep transfer learning using fine-tuning. Models trained from scratch are likely to be more vulnerable.
We will publish the code in an open repository after acceptance of the paper.
@@ Acknowledgments
This work was supported by the Research Council of Finland (Flagship programme: Finnish Center for Artificial Intelligence, FCAI, Grant 356499 and Grant 359111), the Strategic Research Council at the Research Council of Finland (Grant 358247) as well as the European Union (Project 101070617). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the granting authority can be held responsible for them. The authors wish to thank the CSC – IT Center for Science, Finland for supporting this project with computational and data storage resources. We thank Mikko A. Heikkilä and Ossi Räisä for helpful comments and suggestions and John F. Bronskill for helpful discussions regarding few-shot learning.
@@
Appendix A Details of Section 3
@@ A.2 On the assumption of shared scale
In Section 3 we assumed that for LiRA $\sigma_{\mathrm{in}}=\sigma_{\mathrm{out}}$ and $\hat{\sigma}_{\mathrm{in}}=\hat{\sigma}_{\mathrm{out}}$. Using the simplified model formulated in Section 3.4, we show that for large enough number $S$ of examples per class these assumptions are reasonable.
@@ A.3 Proof of Lemma 1
$\begin{cases}t_{\bm{x}}\geq\frac{\hat{\sigma}^{2}\log\tau}{\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}}}+\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2}&\text{ if }\hat{\mu}_{\mathrm{in}}>\hat{\mu}_{\mathrm{out}}\\ t_{\bm{x}}\leq\frac{\hat{\sigma}^{2}\log\tau}{\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}}}+\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2}&\text{ if }\hat{\mu}_{\mathrm{in}}<\hat{\mu}_{\mathrm{out}}.\end{cases}$ (A16)
@@
A.4 Relaxing the assumption of Lemma 1
@@
In Lemma 1 we assume that an attacker has access to the true underlying distribution. However, in real-world settings the precise underlying distribution may not be available for an attacker. In the following, noting that the Equation 8 mainly relies on the true location parameters $\mu_{\mathrm{in}}(\bm{x}),\mu_{\mathrm{out}}(\bm{x})$ and scale parameter $\sigma(\bm{x})$, we relax this assumption of distribution availability so that the attacker only needs an approximated underlying distribution for the optimal LiRA that achieves the performance in Lemma 1.
@@
First, notice that if we completely drop this assumption so that an attacker trains shadow models with an arbitrary underlying distribution, then we may not be able to choose a desired $\textsc{fpr}_{\mathrm{LiRA}}(\bm{x})$. From Equation A22 we have
@@
$\begin{cases}\frac{\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}}}{\hat{\sigma}^{2}}\left(\sigma F_{Z}^{-1}(1-\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))-\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2}+\mu_{\mathrm{out}}\right)\geq 0&\quad\text{ if }\ \hat{\mu}_{\mathrm{in}}>\hat{\mu}_{\mathrm{out}}\\ \frac{\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}}}{\hat{\sigma}^{2}}\left(\sigma F_{Z}^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))-\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2}+\mu_{\mathrm{out}}\right)\geq 0&\quad\text{ if }\ \hat{\mu}_{\mathrm{in}}<\hat{\mu}_{\mathrm{out}}.\end{cases}$ (A31)
Therefore, a sufficient condition about attacker’s knowledge on the underlying distribution for Lemma 1 to hold is
@@
$\begin{cases}\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2}-\mu_{\mathrm{out}}\leq\sigma F_{Z}^{-1}(1-\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))&\quad\text{ if }\ \hat{\mu}_{\mathrm{in}}>\hat{\mu}_{\mathrm{out}}\\ \frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2}-\mu_{\mathrm{out}}\geq\sigma F_{Z}^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))&\quad\text{ if }\ \hat{\mu}_{\mathrm{in}}<\hat{\mu}_{\mathrm{out}}.\end{cases}$ (A33)
@@
Lemma 4 (Lemma 1 with relaxed assumptions).
@@
$\begin{cases}\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2}-\mu_{\mathrm{out}}\leq\sigma F_{Z}^{-1}(1-\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))&\quad\text{ if }\ \hat{\mu}_{\mathrm{in}}>\hat{\mu}_{\mathrm{out}}\\ \frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2}-\mu_{\mathrm{out}}\geq\sigma F_{Z}^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))&\quad\text{ if }\ \hat{\mu}_{\mathrm{in}}<\hat{\mu}_{\mathrm{out}}.\end{cases}$ (A34)

### v5

@@ 4.3 Model to predict dataset vulnerability
The trends seen in Figure˜1 suggest the same power law relationship that we derived for the simplified model of membership inference in Section˜3. We fit a linear regression model to predict $\log(\textsc{tpr}-\textsc{fpr})$ for each $\textsc{fpr}=10^{-k},k=1,\dots,5$ separately using the $\log C$ and $\log S$ as covariates with statsmodels (Seabold and Perktold, 2010). The general form of the model can be found in Equation˜13, where $\beta_{S},\beta_{C}$ and $\beta_{0}$ are the learnable regression parameters.
@@
In Section˜D.2, we propose a variation of the regression model that predicts $\log_{10}(\textsc{tpr})$ instead of $\log_{10}(\textsc{tpr}-\textsc{fpr})$ but this alternative model performs worse on our empirical data and predicts $\textsc{tpr}<\textsc{fpr}$ in the tail when $S$ is very large.
We utilise MIA results of ViT-B (Head) (see Table˜A3) as the training data. Based on the $R^{2}$ (coefficient of determination) score ($R^{2}=0.930$ for the model trained on $\textsc{fpr}=0.001$ data), our model fits the data extremely well. We provide further evidence for other fpr in Figure˜A.3 and Table˜A8 in the Appendix. Figure˜4 shows the parameters of the prediction model fitted to the training data. For larger fpr, the coefficient $\beta_{S}$ is around 0.5-0.5, as our theoretical analysis predicts.
Figure 4: Coefficient values for different fpr when fitting a regression model based on Equation˜13 fitted on data from ViT-B (Head) with LiRA (Table˜A3). The error bars display the 95% confidence intervals based on Student’s t-distribution. Theoretical values in the simplified model is shown by pink dotted lines ($\beta_{S}=0.5$ and $\beta_{C}=0$).
@@
  • R-50 (Head): Figure˜5(a) shows that the regression model is robust to a change of the feature extractor, as it is able to predict the tpr for R-50 (Head) (test $R^{2}=0.790$).
  • R-50 (FiLM): Figure˜5(b) shows that the prediction quality is good for R-50 (FiLM) models. These models are fine-tuned with parameter-efficient FiLM (Perez et al., 2018) layers (See Section˜C.1). Tobaben et al. (2023) demonstrated that FiLM layers are a competitive alternative to training all parameters. We supplement the MIA results of Tobaben et al. (2023) with own FiLM training runs. Refer to Table˜A7 in the Appendix.
  • From-Scratch-Training: Carlini et al. (2022) provide limited results on from-scratch-training. To the best of our knowledge these are the only published LiRA results on image classification models. Figure˜5(b) displays that our prediction model underestimates the vulnerability of the from-scratch trained target models. We have identified two potential explanations for this: 1 In from-scratch-training all weights of the model need to be trained from the sensitive data and thus potentially from-scratch-training could be more vulnerable than fine-tuning. 2 The strongest attack in Carlini et al. (2022) uses data augmentations to improve the performance. We are not using this optimization. Additionally, as noted in Remark˜3, our theoretical analysis suggests that the power-law would not always hold for general from-scratch training.
(a) The dots show the median tpr for the train set (ViT-B; Table˜A3) and the test set (R-50; Table˜A4) over six seeds (datasets: Patch Camelyon, EuroSAT and CIFAR100). The linear model is robust to changing the feature extractor from ViT-B to R-50.
(b) Regression model is robust to changing the fine-tuning method from Head to FiLM, but from scratch training seems to be more vulnerable than predicted. 1 left: fine-tuned with FiLM (see Table˜A7) 2 right: trained from scratch. Data is from Carlini et al. (2022) .
Figure 5: Performance of the regression model based on Equation˜13 fitted on data from Table˜A3.
@@ 4.4 Individual MIA vulnerability
In order to assess the per-sample MIA vulnerability, we run the experiment with $257$ models with each of them once acting as a target and as a shadow model otherwise, compute the tpr at fpr for every sample separately. In Figure˜6, we display the individual vulnerability as a function of $S$ (shots) for Patch Camelyon. The plot shows the maximal vulnerability of all samples and different quantiles over six seeds. The solid line is the median and the errorbars display the min and max over seeds. The quantiles are more robust to extreme outliers and show decreasing trends already at much lower $S$ than the maximum vulnerability.
When fitting the model in Equation˜13 we observe that the coefficients $\beta_{S}$ that model the relationship between vulnerability and examples per class for the quantiles are 0.5603-0.5603, 0.5688-0.5688 and 0.4796-0.4796 which is close to the theoretical value of 0.5-0.5 derived in the theoretical analysis in Section˜3.4. However, the maximum vulnerability decreases with a lower slope of 0.2695-0.2695 which is considerably smaller. With larger $S$ the slope of the max vulnerability increases, e.g., with $S\geq 32768$ the coefficient is 0.3478-0.3478 suggesting that higher $S$ is required for the most vulnerable points.
@@
Table 1: The minimum $S$ with CC=22 predicted by the models in Sections˜4.3 and 4.4 to empirically match the DP bounds ($\delta$$10^{-5}$) calculated through (Kairouz et al., 2015) in terms of tpr at fpr.
$\epsilon$ min $S$ in average case min $S$ in worst-case fpr=$0.1$ fpr=$0.01$ fpr=$0.001$ fpr=$0.1$ $0.25$ 5 4005\,400 69 00069\,000 320 000320\,000 $5.5\times 10^{9}$ $0.50$ 1 1001\,100 16 00016\,000 88 00088\,000 $2.6\times 10^{8}$ $0.75$ $360$ 5 9005\,900 38 00038\,000 $3.5\times 10^{7}$ $1.00$ $160$ 2 7002\,700 19 00019\,000 $7.0\times 10^{6}$
@@ 4.5 Comparison between empirical models and universal DP bounds
Using the translation between the $(\textsc{tpr},\textsc{fpr})$ to $(\epsilon,\delta)$-DP guarantees proposed by Kairouz et al. (2015), we compute the minimum $S$ predicted by our empirical models such that the predicted tpr matches the theoretical bound at target DP level. (See Appendix˜E for a more detailed description.) Table˜1 shows the resulting lower bounds of $S$ for various DP levels and values of fpr. While our empirical observations do not provide any formal guarantees, the comparison serves as an illustration to better understand the different requirements of the average and individual vulnerability. We can see that both for the average and the worst case, obtaining a low fpr would require a large amount of samples per class in order to match the tpr for meaningfully strong DP bounds.
@@ 5 Discussion
Trying to bridge empirical MIA vulnerability and formal DP guarantees is not an easy task because of different threat models and different nature of bounds (statistical vs. universal). While we were able to show both theoretically (Section˜3) and empirically (Section˜4) that having more examples per class provides protection against MIA in fine-tuned neural networks, the numbers required for significant protection (Section˜4.5) limit the practical utility of this observation.
@@
Our formal analysis focuses on a setting that can be linked to fine-tuning. As shown in Figure˜5(b), from-scratch training likely has higher vulnerability. Formally analyzing more models is an interesting area for future research.
Since our MIA evaluation assumes that only a target point is known to the adversary and the rest of the dataset is random, this stochasticity likely introduces some protection. Hence the power-law may be invalidated under stronger MIA settings where the adversary has access to other points in the private dataset (Bai et al., 2025).
Our experiments show that there is a difference between classes in terms of vulnerability and an interesting direction for future work is to understand the properties of classes that influence this vulnerability, e.g., variability within or between classes or their separability. Another direction for future work is understanding the impact of pre-training data on the vulnerability of fine-tuning data.
@@
We mostly consider LiRA in our paper which is optimal for our simplified model of membership inference (Section˜3.4) but for the transfer-learning experiments (Section˜4) there might be stronger attacks in the future. Furthermore, our simplified model assumes well-behaved underlying distributions, meaning that the data is normally distributed around the class centres. We leave the analysis of other data distributions (e.g., heavy-tailed distributions) to future work.. Formal bounds on MIA vulnerability would require something like DP. In addition, both our theoretical and empirical analysis focus on deep transfer learning using fine-tuning. Models trained from scratch are likely to be more vulnerable.
We provide the code in an open repository [^2] https://github.com/DPBayes/impact-dataset-properties-MI-vulnerability-deep-TL. All used pre-trained models and datasets are publicly available.
@@ Acknowledgments
This work was supported by the Research Council of Finland (Flagship programme: Finnish Center for Artificial Intelligence, FCAI, Grant 356499 and Grant 359111), the Strategic Research Council at the Research Council of Finland (Grant 358247), the European Union (Project 101070617) as well as JSPS KAKENHI Grant Number 25KJ1515. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the granting authority can be held responsible for them. he authors wish to acknowledge CSC – IT Center for Science, Finland, for computational resources. We thank Mikko A. Heikkilä and Ossi Räisä for helpful comments and suggestions and John F. Bronskill for helpful discussions regarding few-shot learning.
@@
Contents
Appendix A Details of Section˜3
@@ A.2 On the assumption of shared scale
In Section˜3 we assumed that for LiRA $\sigma_{\mathrm{in}}=\sigma_{\mathrm{out}}$ and $\hat{\sigma}_{\mathrm{in}}=\hat{\sigma}_{\mathrm{out}}$. Using the simplified model formulated in Section˜3.4, we show that for large enough number $S$ of examples per class these assumptions are reasonable.
@@ A.3 Proof of Lemma˜1
[…]
@@
A.4 Offline LiRA
Carlini et al. (2022) proposes offline LiRA that only trains OUT shadow models for computational efficiency. Instead of likelihood ratio, the score for offline LiRA is given as
$\Lambda(\bm{x})=\Pr_{Z}(\hat{\mu}_{\mathrm{out}}(\bm{x})+\hat{\sigma}_{\mathrm{out}}(\bm{x})Z\leq t_{\bm{x}}),$ (A29)
where $Z$ is a location-scale distribution with standard location and unit scale. Then tpr of offline LiRA becomes
$\textsc{tpr}_{\textrm{off}\mathrm{LiRA}}(\bm{x})=\Pr_{\mathcal{D}_{\mathrm{target}}\sim{\mathbb{D}}^{|\mathcal{D}|},\phi^{M}}\left(\Lambda(\bm{x})\geq\tau\mid(\bm{x},y)\in\mathcal{D}_{\mathrm{target}}\right),$ (A30)
where $\gamma$ is a tunable parameter. fpr is also given similarly. Below we show a result for offline LiRA similar to Lemma˜1, but under a slightly more strict assumption that an attacker accurately estimates the IN/OUT score distributions. In the following, as for online LiRA, we assume $\sigma_{\mathrm{in}}=\sigma_{\mathrm{out}}$.
Lemma 5 (Per-example offline LiRA vulnerability).
Suppose that $t_{\bm{x}}$ follows the normal distribution with means $\mu_{\mathrm{in}}(\bm{x}),\mu_{\mathrm{out}}(\bm{x})$ and standard deviation $\sigma(\bm{x})$. Assume that an attacker has access to the underlying distribution $\mathbb{D}$}. Then for a large enough number of examples per class and infinitely many shadow models, the offline LiRA vulnerability of a fixed target example is
$\textsc{tpr}_{\mathrm{offLiRA}}(\bm{x})=\Phi\left(\Phi^{-1}\left(\textsc{fpr}_{\mathrm{offLiRA}}\right)+\frac{\mu_{\mathrm{in}}(\bm{x})-\mu_{\mathrm{out}}(\bm{x})}{\sigma(\bm{x})}\right)$ (A31)
where $\Phi$ is the standard normal cdf.
Proof.
For infinitely many shadow models, we have
$\textsc{fpr}_{\mathrm{offLiRA}}(\bm{x})=\Pr_{\mathcal{D}_{\mathrm{target}}\sim{\mathbb{D}}^{|\mathcal{D}|}}\left(\Lambda(\bm{x})\geq\tau\mid(\bm{x},y)\notin\mathcal{D}_{\mathrm{target}}\right),$ (A32)
and the score is now
$\Lambda(\bm{x})=\Pr_{\eta}(\mu_{\mathrm{out}}(\bm{x})+\sigma(\bm{x})\eta\leq t_{\bm{x}}),$ (A33)
where $\eta\sim\mathcal{N}(0,1)$. When $(\bm{x},y)\notin\mathcal{D}_{\mathrm{target}}$, $t_{\bm{x}}=\mu_{\mathrm{out}}(\bm{x})+\sigma(\bm{x})Z$. Thus we have
$\textsc{fpr}_{\mathrm{offLiRA}}$ $=\Pr_{Z}\left(\Lambda(\bm{x})\geq\gamma\right)$ (A34)
$=\Pr_{Z}\left(\Pr_{\eta}(\mu_{\mathrm{out}}(\bm{x})+\sigma(\bm{x})\eta\leq\mu_{\mathrm{out}}(\bm{x})+\sigma(\bm{x})Z)\geq\gamma\right)$ (A35)
$=\Pr_{Z}\left(\Pr_{\eta}(\eta\leq Z)\geq\gamma\right)$ (A36)
$=\Pr_{Z}(\Phi(Z)\geq\gamma)$ (A37)
$=1-\gamma.$ (A38)
On the other hand, when $(\bm{x},y)\in\mathcal{D}_{\mathrm{target}}$, $t_{\bm{x}}=\mu_{\mathrm{in}}(\bm{x})+\sigma(\bm{x})Z$. Thus we obtain
$\textsc{tpr}_{\mathrm{offLiRA}}$ $=\Pr_{Z}\left(\Pr_{\eta}(\mu_{\mathrm{out}}(\bm{x})+\sigma(\bm{x})\eta\leq\mu_{\mathrm{in}}(\bm{x})+\sigma(\bm{x})Z)\geq\gamma\right)$ (A39)
$=\Pr_{Z}\left(\Pr_{\eta}\left(\eta\leq\frac{\mu_{\mathrm{in}}(\bm{x})-\mu_{\mathrm{out}}(\bm{x})}{\sigma(\bm{x})}+Z\right)\geq 1-\textsc{fpr}_{\mathrm{offLiRA}}\right)$ (A40)
$=\Pr_{Z}\left(\Pr_{\eta}\left(\eta\leq-\frac{\mu_{\mathrm{in}}(\bm{x})-\mu_{\mathrm{out}}(\bm{x})}{\sigma(\bm{x})}-Z\right)\leq\textsc{fpr}_{\mathrm{offLiRA}}\right)$ (A41)
$=\Pr_{Z}\left(-\frac{\mu_{\mathrm{in}}(\bm{x})-\mu_{\mathrm{out}}(\bm{x})}{\sigma(\bm{x})}-Z\leq\Phi^{-1}(\textsc{fpr}_{\mathrm{offLiRA}})\right)$ (A42)
$=\Phi\left(\Phi^{-1}\left(\textsc{fpr}_{\mathrm{offLiRA}}\right)+\frac{\mu_{\mathrm{in}}(\bm{x})-\mu_{\mathrm{out}}(\bm{x})}{\sigma(\bm{x})}\right).$ (A43)
Consequently, the power-law also holds for offline LiRA in the simplified model:
Corollary 6 (Per-example offline LiRA power-law).
Fix a target example $(\bm{x},y)$. For the simplified model with arbitrary $C$ and infinitely many shadow models, the per-example offline LiRA vulnerability is given as
$\log(\textsc{tpr}_{\mathrm{offLiRA}}(\bm{x})-\textsc{fpr}_{\mathrm{offLiRA}}(\bm{x}))\\ =-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{offLiRA}}(\bm{x}))^{2}+\log\frac{|\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle|}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}\sqrt{2\pi}}+\log(1+\xi(S))$ (A44)
$\log(\textsc{tpr}_{\mathrm{offLiRA}}(\bm{x})-\textsc{fpr}_{\mathrm{offLiRA}}(\bm{x}))\approx-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))^{2}+\log\frac{|\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle|}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}\sqrt{2\pi}}.$ (A45)
A.5 Relaxing the assumption of Lemma˜1
@@
In Lemma˜1 we assume that an attacker has access to the true underlying distribution. However, in real-world settings the precise underlying distribution may not be available for an attacker. In the following, noting that the Equation˜8 mainly relies on the true location parameters $\mu_{\mathrm{in}}(\bm{x}),\mu_{\mathrm{out}}(\bm{x})$ and scale parameter $\sigma(\bm{x})$, we relax this assumption of distribution availability so that the attacker only needs an approximated underlying distribution for the optimal LiRA that achieves the performance in Lemma˜1.
@@
First, notice that if we completely drop this assumption so that an attacker trains shadow models with an arbitrary underlying distribution, then we may not be able to choose a desired $\textsc{fpr}_{\mathrm{LiRA}}(\bm{x})$. From Equation˜A22 we have
@@
[…]
Therefore, a sufficient condition about attacker’s knowledge on the underlying distribution for Lemma˜1 to hold is
@@
[…]
@@
Lemma 7 (Lemma˜1 with relaxed assumptions).
@@
[…]